In a school building a classroom contains 60 chairs. The number of chairs per row is three more than thrice the number of rows. Find the number of rows and number of chairs per row.
Answer by ikleyn(52781)

.

Let x be the number of rows.


Then the number of chairs in each row is  3x + 3.


The total chairs equation is


    w*(3w + 3) = 60,   or

    3w*(w+1) = 60

    w*(w+1) = 60/3 = 20,


and then your common sense gives you the unique solution  w = 4.


ANSWER.  There are 4 rows of chairs, with 3*4+3 = 15 chairs in each row.


CHECK.   4*15 = 60 chairs, in total.    ! Correct !

Solved.